# One Stone, Four Birds: A Comprehensive Solution for QA System Using Supervised Contrastive Learning

### v1

@@ IV-B2 Solution for T-2 out-of-domain detection
The solution for T-2 is very simple: we still calculate MDist of $\bm{x}_{test}$ to the nearest intent centroid as T-1, but the distance is used directly as the detection $sco$:
$sco(\bm{x}_{test})={\underset{c\in{\mathcal{C}}}{min}}\ MDist(\bm{h}_{\text{test}},\bar{\bm{h}}_{c})$ (13)
The distance itself can serve as the $sco$ and be an effective detection indicator, since a larger value means that even taking the shortest distance, current example is still far from / has less similarity to any known, IND category centroids. As shown in Figure 2-T-2, this suggests that the current $\bm{x}_{test}$ is more likely to be an unseen, i.e., OOD text.
In practical usage, as introduced in Section III-C, $sco$ is further compared with a threshold $\lambda$, and texts with an exceeding distance will be judged as OOD examples.
@@ IV-B3 Solution for T-3 new intent discovery
To conduct intent discovery, the first step is to obtain the OOD examples to be clustered. Here, we first find a threshold $\lambda$ with a policy that when the model in T-2 gets a true positive rate, TPR=90% in T-2 on validation data (i.e., 90% OOD examples are correctly recognized), then we filter out OOD texts by comparing $sco$ with this $\lambda$.
Next, since SCL has provided a well-structured feature space, we apply KMeans, one of the simplest unsupervised clustering algorithms, to solve T-3. Taking the detected OOD samples $\bm{x}_{j^{\prime}}^{ood}$, KMeans clusters and assigns pseudo-labels $\hat{y}_{j^{\prime}}$ to them. The detailed procedure in algorithm is provided in Appendix A. The pseudo-labeled new intent examples $\{\bm{x}_{j^{\prime}}^{ood},\hat{y}_{j^{\prime}}\}$ will serve as the training data for the final T-4.
@@ V Experiment
V-A Baselines
Fine-tuning, FT (All tasks): Despite being the most common training method, CE-based fine-tuning remains one of the most powerful methods actively used among all tasks [16, 31, 41]. Thus this time we applied it as a baseline of RepL, and tested the effects throughout all four tasks. The downstream detection and clustering method are all the same as SCL.
PTO-Sup [38] (T-2): This work proposed a Prefix Tuning-based OOD detection method, training a generative GPT-2 model on IND data and then using the likelihood as $sco$. It assumes that the unseen OOD text should let model output a lower likelihood. This method originally supports unsupervised learning, but for a fair comparison, we applied its supervised mode with IND labels available.
Layer-LOF [27] (T-2): This paper addressed that pre-trained language models like BERT may over-rationalize the input. That is, even it is an OOD input, the last-layer embedding may tend to be more like IND. Based on this observation, it proposed to consider a decision of multiple layers after fine-tuning, and finally used local outlier factor (LOF) with a voting mechanism for final decision. In this paper, we simply averaged its best-four layer results as the final result.
MTP-KNNCL [44] (T-3): The state-of-the-art NID method that first pre-trains the model with a Masked Language Modeling (MLM) target to obtain prior semantic knowledge, then performs KNN contrastive learning for better cluster aggregation. The downstream clustering method is KMeans.
DPN [45] (T-3): This Decoupled Prototypical Network first initializes some prototypes for known and unknown intents and aligns them as a bipartite matching problem. Unaligned prototypes represent new intents, and all embeddings are further optimized using semantic-aware prototypical learning. The downstream method is also KMeans.
V-B Datasets
We evaluate the methods using three popular intent classification datasets. To measure the performance of T-3 and T-4 quantitatively, we randomly select approximately 75% of categories as the known IND intents, treating the remaining 25% as the unknown, OOD intents to discover new intents from them. Additionally, due to a shortage of test data for T-4, we split around 20% of the training data to create a second test set, Test II.
The original data composition is shown in Table I, with examples from each dataset provided in Table V. Details on IND/OOD intent segmentation and reorganized data composition are shown in Tables II and III, respectively. To prevent confusion due to multiple data partitions and evaluation procedures, a quick reference for data usage is provided in Table IV, with a more detailed explanation in following Section V-D.
CLINC [54]: This is a widely-used dataset in OOD detection. It originally provides 150 intents of texts as the IND data and one OOD intent data (consisting of various abnormal inputs without specific labels). In this experiment, since we cannot calculate clustering metrics on this unlabeled OOD data, we discard it, and randomly select 112 IND intents as the new known IND intents and the remaining 38 intents as the new unknown OOD intents.
BANKING [55]: This intent classification dataset concerns online banking services with 77 intents. We randomly select 57 intents as the IND intents and 20 intents as the OOD intents.
SNIPS [56]: This is another intent classification dataset extracted from a voice assistant. Five out of seven intents are randomly selected to serve as the IND intents and remaining two intents are regarded as the OOD intents.
@@ V-B Datasets
TABLE II: IND, OOD intent segmentation details
Dataset original number of categories num. of selected IND categories num. of selected OOD categories
CLINC 150 112 38
BANKING 77 57 20
SNIPS 7 5 2
TABLE III: Re-organized data composition
Dataset num. of cate. Train Val Test I Test II
CLINC 150 12,000 3,000 4,500 3,000
BANKING 77 7,463 1,000 3,080 1,540
SNIPS 7 12,300 700 700 700
TABLE IV: Data division usage in training/evaluations in this paper
IND part OOD part
Train Pre-training No usage
Val Pre-, T-4 training validation
Test I T-1 test, T-2 test T-2 test, T-3 test
Test II T-4 test
TABLE V: Examples of each dataset
@@
V-C Metrics
We use the following metrics for model evaluation. Basic concepts are provided here, with detailed calculations in Appendix B.
V-C1 For Task 1 and 4: F1 values
Since T-1 and T-4 are both classification problems, we use Micro and Macro F1 scores for evaluation.
V-C2 For Task 2: AUROC, AUPR, FPR90
@@ V-C3 For Task 3: NMI, ARI, ACC
V-D Data usage and training/evaluation order
In this paper, since there are many tasks and data divisions involved in our evaluations, we explain the training and evaluation order here with Figure 3. A complete procedure in algorithm form can also be founded in Appendix D.
As shown in Figure 3, in the Pre-training stage, the model was trained using SCL with Train - IND data. During training, Val - IND/OOD data were used to determine the stopping point, specifically when the OOD detection AUROC on Val set reached its peak.
After pre-training, the basic intent classification performance of the initial model (T-1) was tested on Test I - IND data, while the OOD detection ability (T-2) was tested on Test I - IND/OOD data. The filtered OOD texts (refer to Section IV-B3 for the filtering strategy) were saved and sent to T-3 for new intent discovery. Finally, using the pseudo-labeled old IND data (labeled by initial model) and new “IND” data (labeled in T-3), the initial model was re-trained, and the classification performance was tested on Test II set (T-4).
Figure 3: Diagrams of the evaluation workflow and data division used in each task
V-E Implementation settings
We adopted BERT-base [8] and MPNet-base [57] as the base encoders for both our methods and baselines (expect for PTO-Sup that used GPT-2 small [35]). These encoders were chosen because BERT is the most commonly used pre-trained encoder, while MPNet, a BERT-like variation, enhances context learning ability for better sentence encoding [58]. For SCL setting, we set dropout rate $p\textit{=}0.1$, $n_{\text{views}}\textit{=}2$ for “view” generation, and $\tau\textit{=}0.1$ for calculating the loss. During the pre-training stage, we used the OOD detection AUROC on Val set to evaluate the training process for our SCL and FT methods, saving only the model with the highest $AUROC_{val}$. The training time reported in Table VI also corresponds to the training time of the best model. All experiments were conducted three times, and the results were averaged to obtain the final result. The experiments were run on a machine with an Ubuntu 18.04 operating system, an Nvidia A6000 GPU, an Intel i9-10980X CPU, and 128GB RAM.
VI Results and Analysis
VI-A Training time
TABLE VI: Pre-training time and validation result comparison
@@ VI-A Training time
FT 18 4min10s 95.79 14 3min 93.81 15 3min15s 98.09
SCL (Ours) 7 1min10s 96.37 10 2min30s 94.97 12 2min40s 98.15
TABLE VII: T-1: intent classification result on Test I - IND set
Dataset CLINC BANKING SNIPS
FT-FullyConnect 97.23 97.22 87.32 85.69 98.83 98.74
SCL-NN (Ours) 97.23 97.23 94.17 94.16 98.83 98.75
TABLE VIII: T-2: IND/OOD detection result on Test I - IND/OOD set, the best result is in bold, the runner-up result is underlined
@@
PTO-Sup 81.67 55.01 44.3 40min 79.36 56.44 55.0 72min 97.53 93.87 6.4 3min20s
Layer-CE 95.67 86.71 10.7 78min 92.89 81.67 20.3 47min 93.47 87.34 15.6 18min
FT-NDist 95.64 85.96 9.7 4min10s 92.73 80.12 18.9 3min 96.56 90.43 9.2 3min15s
SCL-NDist (Ours) 96.15 88.41 8.5 1min10s 93.44 82.06 15.7 2min30s 97.41 94.36 4.3 2min40s
Note: All results are calculated under ‘OOD example as class 1 (positive) and IND example as class 0 (negative)’ condition.
TABLE IX: T-3: New intent discovery result on Test I - OOD set, the best result is in bold, the runner-up result is underlined
Dataset CLINC BANKING SNIPS
[…]
MTP-KNNCL 92.36 79.70 83.60 6min10s 83.33 69.75 80.75 6min15s 78.04 85.50 96.26 6min40s
DPN 89.93 75.19 82.89 6min30s 82.51 69.04 76.50 6min 68.85 77.72 94.12 6min30s
FT-KMeans 86.81 67.71 76.27 4min10s 81.88 64.88 73.86 3min 66.16 70.29 91.98 3min15s
SCL-KMeans (Ours) 93.08 81.14 85.67 1min10s 84.44 70.64 81.16 2min30s 70.94 75.82 93.58 2min40s
We first show the model training time by SCL and CE fine-tuning in Table VI. It can be found that our model always reached the stop point earlier with higher Val AUROC scores, paving the way for a better performance in the following downstream tasks.
VI-B Result on task 1: user intent classification
@@
The initial classification results on Test I - IND data are shown in Table VII. Both SCL and FT achieved very high accuracy on CLINC and SNIPS datasets. However, on BANKING, our method showed its strength, achieving significantly better F1 scores on both metrics. This is because, compared with the other two datasets, examples in BANKING have a longer input length while the inquiry expressions are also more diverse (as shown in Table V), which greatly increased classification difficulty. Despite this, our model maintained a high accuracy, demonstrating its robustness even with longer and noisier inputs.
VI-C Result on task 2: out-of-domain detection
@@
The results on T-2 are shown in Table VIII. It can be observed that while FT provided solid results, our method performed better and achieved nearly all the best results. Layer-LOF that applied a layer-voting strategy also obtained slightly better results than FT on the first two datasets. However, the cost is a very long training time by the extensive computations required to measure information redundancy between layers. The random layer selection strategy also increased uncertainty, making it less practical.
@@
On the other hand, PTO-Sup’s generative OOD detection method seemed work badly except on SNIPS. We speculate the reason as SNIPS exhibited more “background-shift” between IND/OOD examples. For instance, there are two intents belonging to IND and OOD parts shared the topic of “watching a movie”, although one was looking for movie information and another involved planning a cinema visit. Distinguishing this kind of OOD texts is a strong point of generative and textual perplexity-based methods but a weak point of CE fine-tune-based methods [36]. This is also evident from the performance drop of the other two fine-tuning-based baselines on SNIPS. Our method, however, still held strong across all situations, showing its superiority on OOD detection.
VI-D Result on task 3: new intent discovery
Table IX shows the results for Task 3 - new intent discovery. The feature optimization capabilities of SCL over CE fine-tuning are proven again by higher clustering scores on all datasets (We will provide the embedding space visualization evidence in following Section VI-F). As for the baselines, MTP-KNNCL also showed good clustering performance especially on SNIPS, where the other methods performed badly. We speculate that this is likely due to the existence of “background shift” discussed in T-2 result, where the unsupervised MLM target and semi-supervised KNN-ConL techniques used by MTP-KNNCL mitigated the risk of “over-training” common in supervised training settings. However, the cost is a slightly longer training time.
VI-E Result on task 4: continual learning
TABLE X: T-4: Continual learning result on Test II set
@@ VI-E Result on task 4: continual learning
Ref-Ori. FT model 98.30 98.31 87.11 85.93 99.40 99.40
Ref-Ori. SCL model 98.39 98.40 93.60 93.63 99.40 99.40

### v2

@@ IV-B2 Solution for T-2 out-of-domain detection
The method to detect OOD examples is also very simple: we still calculate the embedding distance MDist between sample $\bm{x}_{test}$ and the nearest intent centroid as T-1, but the distance is used directly as the detection $sco$:
$sco(\bm{x}_{test})={\underset{c\in{\mathcal{C}}_{kn}}{min}}\ MDist(\bm{h}_{\text{test}},\bar{\bm{h}}_{c})$ (14)
This text representation distance itself can serve as the $sco$ and be an effective detection indicator, since a larger value means that even taking the shortest distance, current example is still far from, or say has less similarity to any known, IND category centroids. As shown in Figure 2-T-2, this suggests that the current $\bm{x}_{test}$ is more likely to be an unseen, i.e., OOD text. In practical usage, as introduced in task definition (Section III-C), this $sco$ is further compared with a threshold $\lambda$, and texts with an exceeding distance will be judged as OOD examples.
@@ IV-B3 Solution for T-3 new intent discovery
To conduct new intent discovery, the first step is to obtain the unknown data to be clustered from T-2. We achieve this by first finding a threshold $\lambda$ with a policy that when the model in T-2 gets a true positive rate, TPR=90% on validation data (i.e., 90% OOD examples in validation set are correctly recognized), then we filter out OOD texts by comparing $sco$ with this $\lambda$.
Next, since SCL is able to provide a well-structured feature space, we do not conduct any further feature pre-processing or analysis but directly apply KMeans, one of the simplest unsupervised clustering method, as T-3’s solution. Taking the representations of detected OOD samples $\bm{x}_{j^{\prime}}^{ood}$, KMeans clusters and assigns pseudo-labels $\hat{y}_{j^{\prime}}$ to them. The detailed clustering procedure of KMeans in algorithm is provided in Appendix A. Finally, the pseudo-labeled new intent examples $\{\bm{x}_{j^{\prime}}^{ood},\hat{y}_{j^{\prime}}\}$ will serve as the training data for the final task T-4.
@@ IV-B4 Solution for T-4 continual learning
This training method can be regarded as a straightforward “replay” method to mitigate forgetting of the old intents. We choose this strategy because the goal of continual learning in our QA system is to enable the model to classify all categories (both old and new intents) simultaneously, rather than in separate stages like other studies. Additionally, since the old intent data are readily available and can be stored without significant cost, as a result, the replay strategy is simple, easy-to-deploy and well-suited to our objectives. Experiment results (Section VI-D) also demonstrate that this replay policy effectively prevents forgetting, with only minimal performance degradation observed in classifying old data.
@@ V Experiment
V-A Datasets, division, and evaluation order
To clearly present the evaluation process for the four tasks, we begin by introducing the original dataset, then specify the data divisions applied for the experiment, and finally outline the evaluation workflow based on these divisions.
V-A1 Dataset
We use three popular public intent classification datasets for the experiment.
CLINC [54]: This is a widely-used dataset in intent classification and OOD detection tasks. It provides 150 intents of texts and one OOD intent data (that is, on OOD detection task, the former 150 intents can be treated as the IND intents). Because its OOD part does not contain specific labels that are necessary for T-3 and T-4 evaluation, we only use its 150-intent part data this time.
BANKING [55]: This is an intent classification dataset concerning online banking services, which contains around 13,000 records in 77 intents.
SNIPS [56]: This is another intent classification dataset extracted from a voice assistant, consisting of around 14,000 records in 7 intents.
The original data composition of above datasets is shown in Table I. Some examples are also provided in Table II. Since all data are well-formed, which keep a relatively good data balance and do not contain personal information, we did not take further pre-processing steps but used them as they are.
@@ V-A1 Dataset
TABLE II: Examples of each dataset
@@
V-A2 Data division for the experiment
We apply two key divisions to adapt the data for the experimental requirements. The first is the IND/OOD intent split, where approximately 75% number of the original intents are randomly selected as the known IND intents, with the remaining 25% treated as the unknown OOD intents. Specifically, in CLINC, BANKING and SNIPS datasets, 112, 57, and 5 intents were randomly selected as the known IND intents, respectively, while the remaining 38, 20, and 2 intents were designated as the unknown OOD intents.
Additionally, since T-4 requires the recognition results from T-2 and T-3 for the final re-training (refer to Section III-E for this process and Figure 3 for workflow), the original single test set is insufficient to evaluate all four tasks, especially as it may cause data leakage for T-4. To address this, we further split around 20% data from the original training set to create a secondary test set, Test II. The new composition of the re-divided data is presented in Table III.
TABLE III: Re-organized data composition
Dataset IND/OOD cate. Train Val Test I Test II
CLINC 112/38 12,000 3,000 4,500 3,000
BANKING 57/20 7,463 1,000 3,080 1,540
SNIPS 5/2 12,300 700 700 700
V-A3 Evaluation order and data usage
Here we introduce the training and evaluation processes using above divided data. A simple workflow diagram is shown in Figure 3, while the complete procedure in pseudo-code form can also be found in Appendix D.
Training stage: As shown in Figure 3, in the pre-training stage, the model is trained with an SCL target using only the IND part of training data (Train-IND). During training, Val - IND/OOD data are used to determine the stopping point, specifically when the OOD detection AUROC (refer to Section V-B for definition) on Val set reached its peak.
Evaluation stage: After pre-training, the intent classification performance (T-1) of the initial model is first tested on Test I-IND data. The OOD detection ability (T-2) is tested on Test I-IND/OOD data. After that, the filtered OOD texts from T-2 (refer to Section IV-B3 for filtering strategy) are used for new intent discovery (T-3) and the accuracy is then evaluated. Finally, using the pseudo-labeled old IND data (labeled by initial model) and new “IND” data (labeled in T-3), the initial model is re-trained (T-4), and the classification performance is tested on Test II set.
Figure 3: Diagram of the training and evaluation workflow
V-B Metrics
We use the following metrics for model evaluation in each task. Basic concepts are provided here, with detailed calculation processes in Appendix B.
V-B1 For Task 1 and 4: F1 values
Since T-1 and T-4 both belong to classification problems, we use Micro and Macro F1 scores for the evaluation.
V-B2 For Task 2: AUROC, AUPR, FPR90
@@ V-B3 For Task 3: NMI, ARI, ACC
V-C Baselines
Fine-tuning, FT (For all four tasks): Despite being the most common training method, CE-based fine-tuning remains one of the most powerful methods actively used among all tasks [16, 31, 41]. Thus this time we applied it as a baseline of RepL, and tested the effects throughout all four tasks. After fine-tuning, the downstream detection and clustering methods are all the same as SCL.
PTO-Sup [38] (For T-2): This work proposed a Prefix Tuning-based OOD detection method, training a generative GPT-2 model on IND data and then using the likelihood as $sco$. It assumes that the unseen OOD text should let model output a lower likelihood. This method originally supports unsupervised learning, but for a fair comparison, we applied its supervised mode with IND text labels available.
Layer-LOF [27] (For T-2): This paper addressed that pre-trained language models like BERT may over-rationalize the input. That is, even it is an OOD input, the last-layer embedding may tend to be more like IND. Based on this observation, it proposed to consider a decision of multiple layers after fine-tuning, and finally used local outlier factor (LOF) with a voting mechanism for the final decision. In this paper, we simply averaged its best-four layer results as the final result.
MTP-KNNCL [44] (For T-3): The state-of-the-art NID method that first pre-trains the model with a Masked Language Modeling (MLM) target to obtain prior semantic knowledge, then performs KNN contrastive learning for better cluster aggregation. The downstream clustering method is KMeans.
DPN [45] (For T-3): This Decoupled Prototypical Network first initializes some prototypes for known and unknown intents and aligns them as a bipartite matching problem. Unaligned prototypes represent new intents, and all embeddings are further optimized using semantic-aware prototypical learning. The downstream method is also KMeans.
TABLE IV: T-1: intent classification result on Test I - IND set
Dataset CLINC BANKING SNIPS
Method/Metric Micro F1$\uparrow$ Macro F1$\uparrow$ Time/Epoch Micro F1$\uparrow$ Macro F1$\uparrow$ Time/Epoch Micro F1$\uparrow$ Macro F1$\uparrow$ Time/Epoch
FT-FullyConnect 97.23 97.22 4min10s/18 87.32 85.69 3min/14 98.83 98.74 3min15s/15
SCL-NN (Ours) 97.23 97.23 1min10s/7 94.17 94.16 2min30s/10 98.83 98.75 2min40s/12
TABLE V: T-2: IND/OOD detection result on Test I - IND/OOD set, the best result is in bold, the runner-up result is underlined
@@ V-C Baselines
PTO-Sup 81.67 55.01 44.3 40min/8 79.36 56.44 55.0 72min/16 97.53 93.87 6.4 3min20s/13
Layer-LOF 95.67 86.71 10.7 78min/19 92.89 81.67 20.3 47min/12 93.47 87.34 15.6 28min/8
FT-NDist 95.64 85.96 9.7 4min10s/18 92.73 80.12 18.9 3min/14 96.56 90.43 9.2 3min15s/15
SCL-NDist (Ours) 96.15 88.41 8.5 1min10s/7 93.44 82.06 15.7 2min30s/10 97.41 94.36 4.3 2min40s/12
Note: All results are calculated under ‘OOD example as class 1 (positive) and IND example as class 0 (negative)’ condition to highlight the model’s detection ability.
TABLE VI: T-3: New intent discovery result on Test I - OOD set, the best result is in bold, the runner-up result is underlined
@@
MTP-KNNCL 92.36 79.70 83.60 6min10s/20 83.33 69.75 80.75 6min15s/20 78.04 85.50 96.26 6min40s/20
DPN 89.93 75.19 82.89 6min30s/20 82.51 69.04 76.50 6min/20 68.85 77.72 94.12 6min30s/20
FT-KMeans 86.81 67.71 76.27 4min10s/8 81.88 64.88 73.86 3min/14 66.16 70.29 91.98 3min15s/15
SCL-KMeans (Ours) 93.08 81.14 85.67 1min10s/7 84.44 70.64 81.16 2min30s/10 70.94 75.82 93.58 2min40s/12
V-D Other implementation details
We adopted BERT-base [8] and MPNet-base [57] as the base encoders for both our methods and baselines (expect for PTO-Sup that used GPT-2 small [35]). These encoders were chosen because BERT is the most commonly used pre-trained encoder, while MPNet, a BERT-like variation, enhances context learning ability for better sentence encoding [58]. For SCL setting, we set dropout rate $p\textit{=}0.1$, $n_{\text{views}}\textit{=}2$ for “view” generation, and $\tau\textit{=}0.1$ for calculating the loss. During the pre-training stage, we used the OOD detection AUROC on Val set to evaluate the training process for our SCL and FT methods, saving only the model with the highest $AUROC_{val}$. The maximum training epoch was set to 20 for both our and all baseline methods. All experiments were conducted three times, and the results were averaged to obtain the final result. All experiments were running on the same machine with an Ubuntu 18.04 operating system, an Nvidia A6000 GPU, an Intel i9-10980X CPU, and 128GB RAM.
VI Results and Analysis
VI-A Result on task 1: user intent classification
@@
The intent classification results on Test I-IND data are shown in Table IV. We also provide the training time and epoch as an efficiency reference. It can be found that both proposed SCL and traditional FT method achieved very high F1 scores on CLINC and SNIPS datasets. However, things changed on BANKING, where FT method saw a considerable accuracy drop in both metrics, while our method kept a relatively good performance. We analyze that this is because, compared with the other two datasets, the user inquires in BANKING have a longer input length in average, and the inquiry expressions also become more diverse (that is, more detailed and have many patterns, as the examples shown in Table II), which greatly increased classification difficulty. Despite this, the proposed SCL-NN method was still able to give a more inter-class separated and intra-class compact embedding space (a visual prove will be given in Section VII-A), guaranteeing a high accuracy even with longer and noisier inputs.
VI-B Result on task 2: out-of-domain detection
@@
The results evaluating the detection ability of unknown OOD inputs are shown in Table V. Firstly, comparing FT and SCL, it can be observed that while FT provided very solid results, the proposed method performed even better and achieved nearly all the best results. As for the other baselines, Layer-LOF that applied a layer-voting strategy obtained only slightly better results than FT on the first two datasets. Nevertheless, the cost is a much longer training time by the extensive computations required to measure the information redundancy between layers. The random layer selection strategy also increased uncertainty, making it less practical. On the other hand, the generative model-based method PTO-Sup worked badly except on SNIPS, which even obtained the highest AUROC on it. We speculate the reason of this abnormal situation as compared with the other two datasets, the IND/OOD examples in SNIPS exhibited more “background-shift” than usual “semantic-shift”, which can be exactly the strong point of generation-based methods.
@@
Talking more specifically, for instance, there are two intents belonging to IND and OOD intents shared the topic of “watching a movie” in SNIPS, although one was looking for movie information and another involved planning a cinema visit, which is a typical “background-shift”. (On the contrary, “watching a movie” and “read a book” intent pair is totally different in topic and thus is a “semantic-shift”.) This difference is easier for generative and textual perplexity-based methods like PTO-Sup to capture and distinguish, while it can be a weak point of CE fine-tune-based methods [36]. This is also evident from the performance drop of the other two fine-tuning-based baselines on SNIPS. However, it is clear that SCL still held strong on not only the “semantic-shift”, but also the “background-shift” situation, showing its superiority on the OOD intent detection task.
TABLE VII: T-4: Continual learning result on Test II set
@@
Ref.-Initial FT model 98.30 98.31 87.11 85.93 99.40 99.40
Ref.-Initial SCL model 98.39 98.40 93.60 93.63 99.40 99.40
@@
VI-C Result on task 3: new intent discovery
Table VI shows the results for new intent discovery. Comparing SCL and traditional FT method, the advantage of SCL concerning feature optimization capability is proven again by the surpassing clustering scores on all datasets. As for the other baselines, DPN also achieved better scores than FT, but the improvement was not significant. MTP-KNNCL showed very strong clustering performance, achieving second best results to our method on the first two datasets and the best results on SNIPS that other methods performed much worse.
We consider the possible reason as the existence of “background-shift” in SNIPS discussed in T-2 result, where the supervised training style of the other three methods might exacerbate the confusion between those intents. On the other hand, the two techniques, unsupervised MLM target and semi-supervised KNN-ConL, adopted by MTP-KNNCL mitigated the risk of “over-training” in supervised settings, and thus giving a better result. However, the cost is a slightly longer training time.
VI-D Result on task 4: continual learning
Finally, we present the intent classification results after the incremental learning in Table VII. From an overall view (results on all classes), SCL has continued to achieve the best results, and even widened the accuracy gap with FT on all datasets from 0 - 8.5% (on T-1 initial classification) to 4.3% - 10.1%, which we consider as a result of the excellent clustering performance on T-3 and the feature optimization ability of SCL. The concerning forgetting phenomenon also did not appear, as only a slight accuracy decrease on the old IND intent classification was observed after the continual learning. This proves that the adopted “replay” strategy using both pseudo-labeled IND and OOD texts is effective.
In addition to the continuous high accuracy, another advantage of SCL-based re-training is also worth noting. When conducting FT-based re-training, the last fully-connected layer had to be replaced and reinitialized every time to accommodate the new intent number, and it required at least 4-5 epochs to obtain desirable results in our experiments. On the other hand, since SCL focuses on the feature optimization and the label judgment is based on the nearest neighbor algorithm, it can directly use the new data with an extended label set for model re-training in a shorter time. This easy-to-use expandability can save a considerable amount of resources and is meaningful in real-world deployments.